# Identifying Weight-Variant Latent Causal Models

### v6

@@ 4 Identifiable Causal Representations with Weight-variant Linear Gaussian Models
Intuitively, allowing causal influences among latent causal variables to change leads to varying causal influences. Such variations cannot be ’absorbed’ by an invariant nonlinear mapping from $\mathbf{z}$ to $\mathbf{x}$, which breaks the transitivity mentioned in Section 3.2, resulting in identifiability. For example, assume that $z_{1}:=n_{1}$, $z_{2}:=\lambda(\mathbf{u})z_{1}+n_{2}$ and $\mathbf{x}:=\mathbf{f}(z_{1},z_{2})$. If we allow the causal influence $\lambda(\mathbf{u})z_{1}$ to be ’absorbed’ by the mapping $\mathbf{f}$, we would have $z_{1}:=n_{1}$, $z_{2}:=n_{2}$, and the resulting mapping, $\mathbf{f^{\prime}}=\mathbf{f}\circ\mathbf{g}$ where $\mathbf{g}(z_{1},z_{2})=[z_{1},\lambda(\mathbf{u})z_{1}+z_{2}]$, would need to include the original $\mathbf{f}$ as well as the term $\lambda(\mathbf{u})$ in $\mathbf{g}$, to obtain the same obervation $\mathbf{x}$. This means that the resulting mapping $\mathbf{f^{\prime}}$ would change across $\mathbf{u}$ due to the term $\lambda(\mathbf{u})$, which violates the fact that the mapping from $\mathbf{z}$ to $\mathbf{x}$ remains unchanged across $\lambda(\mathbf{u})$ as defined in Eq. (3).
@@ 5 Change of Part of Weights and Partial Identifiability Result
    if $z_{i}$ is a root node or all weights to $z_{i}$, e.g., $\boldsymbol{\lambda}_{i}$ in Eq. (2), meet assumption (v) in Theorem 1, then the true $z_{i}$ is related to the recovered one $\hat{z}_{j}$, obtained by matching the true marginal data distribution $p(\mathbf{x}|\mathbf{u})$, by the following relationship: $z_{i}=s\hat{z}_{j}+c$, where $s$ denotes scaling, $c$ denotes a constant,
@@
    if there exists an unchanged weight from the parent node $z_{j}$ to $z_{i}$, e.g., in ${\lambda}_{j,i}$, in Eq. (2), then $z_{i}$ is unidentifiable,
@@ 6 Learning Causal Representations with Weight-Variant Linear Models
According to the causal model for the latent causal variables, e.g., Eq. (1) and Eq. (2), we construct a linear Gaussian prior with weights and noise depending on $\mathbf{u}$. Since we can pre-define the causal order as mentioned above, let us assume the causal order is $z_{1},z_{2},..z_{n}$. Given this, we can create a fully connected causal graph based on the predefined causal order, where each $z_{i}$ has parent nodes $z_{i}^{\prime}$ for $i^{\prime}<i$. Summarizing the above, we propose the following prior model:
@@
In our implementation, we use one-hot encoding for the additional observed variable $\mathbf{u}$. To enforce assumption (v), we also create new data points with $\mathbf{u}=\mathbf{0}$, which correspond to an output of 00. This allows us to translate assumption (v) into an implemented constraint, such as applying the L1 norm on $\lambda_{i,j}$, as discussed in the last two terms of Eq. 13, where $\gamma_{1}$ and $\gamma_{2}$ are two hyper-parameters used to balance the constraints and the ELBO. Figure 5 depicts the implementation of the proposed method to learn causal representations.
@@
Beyond standard VAEs, recent advanced works aim to explore hierarchical VAEs with the hope that they could improve the evidence lower bound in standard VAE and decrease reconstruction error compared to standard VAEs. More importantly, the stack of latent variables in hierarchical VAEs might learn a feature hierarchy, where different layers capture information at varying levels of detail and abstraction (Kingma et al., 2016; Sønderby et al., 2016; Maaløe et al., 2019; Vahdat and Kautz, 2020). To this end, hierarchical VAE models often employ the following generative model: $z_{1}\rightarrow z_{2}\rightarrow...\rightarrow z_{n}\rightarrow\mathbf{x}$. Interestingly, hierarchical VAE models can be regarded as a special case of the proposed prior model in Equation (10), where the fully-connected graph is reduced to the simpler connections used in the generative model of hierarchical VAEs. Essentially, the proposed SuaVE also aims to learn hierarchical latent variables. Due to permutation indeterminacy, we can model the prior or posterior as hierarchical structures. The difference is that the edges among latent variables of the proposed SuaVE are learned adaptively from data, unlike the simple connections in hierarchical VAEs. Importantly, the proposed SuaVE has rigorous theoretical justification for learning hierarchical latent variables and the structures among them, whereas hierarchical VAEs lack such support without further assumptions. Further exploring the relationship between hierarchical VAEs and latent causal models is both interesting and worthwhile.
@@ 7.1 Synthetic Data
We first conduct experiments on synthetic data, generated by the following process: we divide the latent noise variables into $M$ segments, where each segment corresponds to one conditional variable $\mathbf{u}$ as the segment label. Within each segment, we first sample the mean ${\beta}_{i,1}$ and variance ${\beta}_{i,2}$ from uniform distributions $[-2,2]$ and $[0.01,3]$, respectively. We then sample the weights $\boldsymbol{\lambda}_{i,j}$ from uniform distributions $[0.1,2]$. Then for each segment, we generate the latent causal samples according to the generative model in Eq.3. Finally, we obtain the observed data samples $\mathbf{x}$ by an invertible nonlinear mapping. More details can be found in Appendix.
@@
Figure 8: Performances of the proposed SuaVE in comparison to iVAE, $\beta$-VAE, VAE and CausalVAE in recovering the latent causal variables on synthetic data with different numbers of the latent causal variables. Implementation details can be found in Appendix A.6.
@@
We compared the performance of the proposed SuaVE to some variants of VAE mentioned above. We used the same network architecture for encoder and decoder parts in all these models. In particular, we add a sub-network to model the conditional prior in both iVAE and the proposed SuaVE. We further assign a linear SCM sub-network to model the relations among latent causal variables in the proposed SuaVE. We trained these 5 models on the dataset described above, with different numbers of segments and different numbers of latent causal variables. For each method, we use 5 different random seeds for data sampling. Figure 7 shows the performance on two latent causal variables with different numbers of segment. The proposed SuaVE obtains the score 0.96 approximately for all the different numbers of segment. In contrast, $\beta$-VAE, VAE and CausalVAE fail to achieve a good estimation of the true latent variables, since they are not identifiable. iVAE obtains unsatisfactory results, since its theory only holds for i.i.d. latent variables. Figure 8 shows the performance in recovering latent causal variables on synthetic data with different numbers of the latent causal variables. Figure 9 shows the structural Hamming distance (SHD), a standard metric for comparing graphs via their adjacency matrices, for the recovered skeletons (the undirected version of the graph with all directed edges replaced by undirected edges), causal graphs, and completed partial directed acyclic graphs (CPDAGs, which capture the set of all DAGs that are Markov equivalent), by the proposed SuaVE and CausalVAE. Since CausalVAE is not identifiable, it obtains fully connected graphs, the SHD of the recovered CPDAG is the same as one of the recovered skeletons.
@@ 7.1.1 Experiments on Changes of Part of Weights
In this part, we conduct experiments on changes of part of weighs, to verify the results in Theorem 4. To this end, we consider two cases on the following graph: $z_{1}\rightarrow z_{2}\rightarrow z_{3}\rightarrow z4$. The data details are similar to those of the 4-dimensional case mentioned in Appendix A.6.2, except that $z_{3}$ is a parent node of $z_{4}$. Case 1: the weight on the edge of $z_{1}\rightarrow z_{2}$ (and accordingly, $z_{2}\rightarrow z_{3}$, $z_{3}\rightarrow z_{4}$) remains unchanged (e.g., a constant) across $\mathbf{u}$, while the weights on the remaining edges change across $\mathbf{u}$, Case 2: the weight on the edge of $z_{1}\rightarrow z_{2}$ (and accordingly, $z_{2}\rightarrow z_{3}$, $z_{3}\rightarrow z_{4}$) includes a constant part (e.g., $\lambda_{1,2}(\mathbf{u})+b$, $b$ is a non-zero constant.) across $\mathbf{u}$, while the weights on the remaining edges change across $\mathbf{u}$. Again, the network architecture, optimization, and hyper-parameters are the same as used for experiments of Figure 7 (a). Figures 10 and 11 show the performance of the proposed method in the settings corresponding to Case 1 and Case 2 mentioned above. According to MPC, we observe that unchanged weights in Case 1 or weights involving a constant term in Case 2 lead to non-identifiability. Conversely, changing weights contribute to the identifiability of the corresponding nodes. These empirical results align with the partial identifiability results in Theorem 4.
@@
Figure 11: Case 2: Performances of the proposed method with the change of part of weights. The ground truth of the causal graph is $z_{1}\rightarrow z_{2}\rightarrow z_{3}\rightarrow z_{4}$. From top to bottom: weight on $z_{1}\rightarrow z_{2}$ (and accordingly, $z_{2}\rightarrow z_{3}$, and $z_{3}\rightarrow z_{4}$) includes a non-zero constant part and changing part, e.g., $\lambda(\mathbf{u})+b$. Those results are consistent with the analysis of partial identifiability results in Theorem 4.
@@
Our work is closely related to nonlinear ICA, and our assumptions include the assumptions for identifying the Gaussian noise variables. To verify this point, we conduct experiments on the case where any two latent causal variables $z_{i}$ are independent given $\mathbf{u}$. Data details are the same as mentioned in section A.6.2, but we here enforce $\lambda_{i,j}(\mathbf{u})=0$ for each $i,j$. Again, the network architecture, optimization and hyper-parameters are the same as section A.6.2, except for learning rate (here we set $21r=1e-$ 2). Figure 12 shows the performances of the proposed SuaVE and iVAE. We can see that iVAE is slightly better than the proposed SuaVE, both obtain satisfying results in terms of MPC.
@@ 7.1.3 Performance for violated assumption on distribution
Table 1: Performance for violated assumptions. Uniform (Laplace, Gamma) Noise: the distribution of latent noise variables. Unchanged Weights: weights keep unchanged across $\mathbf{u}$. Violated assumption (v): $\lambda_{i,j}({\mathbf{0}})=c$ where $c$ a non-zero constant. Matching Assumptions: a setting that matches our theoretical assumptions.
@@ 7.2 Image Data
We further verify the proposed identifiability results and method on images from the chemistry dataset proposed in (Ke et al., 2021), which is corresponding to simple chemical reactions where the state of an element can cause changes to another variable’s state. The environment consists of a number of objects whose positions are kept fixed and thus, uniquely identifiable, while the colors (states) of the objects affected by this variable (according to the causal graph) can change. We use a weight-variant linear Gaussian model as mentioned in section A.6.2 for a 3-dimensional case, to generate the latent variables, and obtain the corresponding observational images. A visualization of the environment can be found in Figure 13.
@@
All these observations are consistent with the ground truth latent causal structure, e.g., $z_{1}\rightarrow z_{2}\rightarrow z_{3}$, together with results in Figure 14, demonstrating that the proposed method successfully learns latent causal structures. Similar, we inversigate the change in reconstructions while traversing each learned latent variable obtained by $\beta$-VAE (Higgins et al., 2017) and CausalVAE (Yang et al., 2021). Figure 16 and 17 show traversal results on the learned representations by $\beta$-VAE (Higgins et al., 2017) and CausalVAE (Yang et al., 2021), respectively. we observed that changing each learned variable results in changes to the colors of all objects. Together with the results in Figure 14, these suggests that $\beta$-VAE (Higgins et al., 2017) and CausalVAE (Yang et al., 2021) tend to learn representations that entangle the true underlying latent causal variables and fail to recover the true underlying latent causal structure.
@@
Following Ghassami et al. (2018), we applied the proposed method to fMRI hippocampus dataset (Laumann and Poldrack, 2015), which contains signals from six separate brain regions: perirhinal cortex (PRC), parahippocampal cortex (PHC), entorhinal cortex (ERC), subiculum (Sub), CA1, and CA3/Dentate Gyrus (DG) in the resting states on the same person in 84 successive days. Each day is considered as different $\mathbf{u}$, thus $\mathbf{u}$ is a 84-dimensional vector. Since we are interested in recovering latent causal variables, we treat the six signals as latent causal variables by applying a random nonlinear mapping on them to obtain observed data. We then apply various methods on the observed data to recover the six signals. Figure 19 (a) shows the performance of the proposed SuaVE in comparison to iVAE, $\beta$-VAE, VAE and CausalVAE in recovering the latent six signals. $\beta$-VAE aims to recover independent latent variables, and it obtains an interesting result: enforcing independence (e.g., $\beta=25,50$) leads to worse MPC, and relaxing it (though contracting to its own independence assumption) improves the result (e.g. $\beta=4$). This is because the latent variables given the time index are not independent in this dataset.
@@
Figure 20 (a) show the edges in the anatomical ground truth reported in (Ghassami et al., 2018). Figure 20 (b) show the latent recovered edges by the proposed SuaVE, in the setting where we treat fMRI data as latent causal variables. For comparison, Figure 20 (c) shows the recovered causal graph reported in (Ghassami et al., 2018), in the setting where fMRI data are treated as observed variables. That is, the setting in (b) is more challenging than the setting in (c).

### v7

@@ 4 Identifiable Causal Representations with Weight-variant Linear Gaussian Models
Intuitively, allowing causal influences among latent causal variables to change leads to varying causal influences. Such variations cannot be ’absorbed’ by an invariant nonlinear mapping from $\mathbf{z}$ to $\mathbf{x}$, which breaks the transitivity mentioned in Section 3.2, resulting in identifiability. For example, assume that $z_{1}:=n_{1}$, $z_{2}:=\lambda(\mathbf{u})z_{1}+n_{2}$ and $\mathbf{x}:=\mathbf{f}(z_{1},z_{2})$. If we allow the causal influence $\lambda(\mathbf{u})z_{1}$ to be ’absorbed’ by the mapping $\mathbf{f}$, we would have $z_{1}:=n_{1}$, $z_{2}:=n_{2}$, and the resulting mapping, $\mathbf{f^{\prime}}=\mathbf{f}\circ\mathbf{g}$ where $\mathbf{g}(z_{1},z_{2})=[z_{1},\lambda(\mathbf{u})z_{1}+z_{2}]$, would need to include the original $\mathbf{f}$ as well as the term $\lambda(\mathbf{u})$ in $\mathbf{g}$, to obtain the same obervation $\mathbf{x}$. This means that the resulting mapping $\mathbf{f^{\prime}}$ would change across $\mathbf{u}$ due to the term $\lambda(\mathbf{u})$, which violates the fact that the mapping from $\mathbf{z}$ to $\mathbf{x}$ remains unchanged across $\mathbf{u}$ as defined in Eq. (3).
@@ 5 Change of Part of Weights and Partial Identifiability Result
    if $z_{i}$ is a root node or all weights to $z_{i}$, $\boldsymbol{\lambda}_{i}$ in Eq. (2), meet assumption (v) in Theorem 1, then the true $z_{i}$ is related to the recovered one $\hat{z}_{j}$, obtained by matching the true marginal data distribution $p(\mathbf{x}|\mathbf{u})$, by the following relationship: $z_{i}=s\hat{z}_{j}+c$, where $s$ denotes scaling, $c$ denotes a constant,
@@
    if there exists an unchanged weight ${\lambda}_{j,i}$ from the parent node $z_{j}$ to $z_{i}$, in Eq. (2), then $z_{i}$ is unidentifiable,
@@
Discussion 3: Potential Implications
While partial identifiability may be more realistic in practice, since only some causal influences vary across environments, it still poses challenges in determining whether the latent causal variables have been successfully identified or are practically useful, especially in the absence of ground-truth latent variables. Nevertheless, such partial identifiability results may be valuable in downstream applications, particularly in domain shift scenarios such as domain adaptation, domain generalization, and transfer learning (Kong et al., 2022; Li et al., 2023; Liu et al., 2025). In these settings, partially identified latent causal variables that exhibit stable causal relationships with the target label across environments may capture invariant mechanisms useful for robust prediction. Their relevance can be empirically assessed through predictive performance, stability under distribution shifts.
@@ 6 Learning Causal Representations with Weight-Variant Linear Models
According to the causal model for the latent causal variables, i.e., Eq. (1) and Eq. (2), we construct a linear Gaussian prior with weights and noise depending on $\mathbf{u}$. Since we can pre-define the causal order as mentioned above, let us assume the causal order is $z_{1},z_{2},..z_{n}$. Given this, we can create a fully connected causal graph based on the predefined causal order, where each $z_{i}$ has parent nodes $z_{i}^{\prime}$ for $i^{\prime}<i$. Summarizing the above, we propose the following prior model:
@@
In our implementation, the observed variable $\mathbf{u}$ is represented using one-hot encoding, so that in the experiments $\mathbf{u}$ takes values in a finite discrete set. In this discrete setting, Assumption (v) can in principle be satisfied exactly by fixing $\lambda_{i,j}(\mathbf{u}=0)=0$ by construction and learning $\lambda_{i,j}(\mathbf{u})$ only for $\mathbf{u}\neq 0$. In our implementation, we instead adopt a simple penalty that encourages $\lambda_{i,j}(\mathbf{u}=0)$ to be close to zero, which serves as a convenient implementation choice. This penalty is included in the last two terms of Eq. (13), where $\gamma_{1}$ and $\gamma_{2}$ are hyperparameters that balance the penalty with the ELBO. Figure 5 illustrates the overall implementation of the proposed method.
@@
Beyond standard VAEs, recent advanced works aim to explore hierarchical VAEs with the hope that they could improve the evidence lower bound in standard VAE and decrease reconstruction error compared to standard VAEs. More importantly, the stack of latent variables in hierarchical VAEs might learn a feature hierarchy, where different layers capture information at varying levels of detail and abstraction (Kingma et al., 2016; Sønderby et al., 2016; Maaløe et al., 2019; Vahdat and Kautz, 2020). To this end, hierarchical VAE models often employ the following generative model: $z_{1}\rightarrow z_{2}\rightarrow...\rightarrow z_{n}\rightarrow\mathbf{x}$. Interestingly, hierarchical VAE models can be regarded as a special case of the proposed prior model in Eq. (10), where the fully-connected graph is reduced to the simpler connections used in the generative model of hierarchical VAEs. Essentially, the proposed SuaVE also aims to learn hierarchical latent variables. Due to permutation indeterminacy, we can model the prior or posterior as hierarchical structures. The difference is that the edges among latent variables of the proposed SuaVE are learned adaptively from data, unlike the simple connections in hierarchical VAEs. Importantly, the proposed SuaVE has rigorous theoretical justification for learning hierarchical latent variables and the structures among them, whereas hierarchical VAEs lack such support without further assumptions. Further exploring the relationship between hierarchical VAEs and latent causal models is both interesting and worthwhile.
@@ 7.1 Synthetic Data
We first conduct experiments on synthetic data, generated by the following process: we divide the latent noise variables into $M$ segments, where each segment corresponds to one conditional variable $\mathbf{u}$ as the segment label. Within each segment, we first sample the mean ${\beta}_{i,1}$ and variance ${\beta}_{i,2}$ from uniform distributions [2,2][-2,2] and $[0.01,3]$, respectively. Subsequently, we sample the weights ${\lambda}_{i,j}$ from a uniform distribution over $[0.1,2]$. Then for each segment, we generate the latent causal samples according to the generative model in Eq. (2). Finally, we obtain the observed data samples $\mathbf{x}$ by an invertible nonlinear mapping. More details can be found in Appendix.
@@
Figure 8: Performances of the proposed SuaVE in comparison to iVAE, $\beta$-VAE, VAE and CausalVAE in recovering the latent causal variables on synthetic data with different numbers of the latent causal variables. Details can be found in Appendix A.6.
@@
We compared the performance of the proposed SuaVE to some variants of VAE mentioned above. We used the same network architecture for encoder and decoder parts in all these models. In particular, we add a sub-network to model the conditional prior in both iVAE and the proposed SuaVE. We further assign a linear SCM sub-network to model the relations among latent causal variables in the proposed SuaVE. We trained these 5 models on the dataset described above, with different numbers of segments and different numbers of latent causal variables. For each method, we use 5 different random seeds for data sampling. Figure 7 shows the performance on two latent causal variables with different numbers of segment. The proposed SuaVE obtains the score 0.96 approximately for all the different numbers of segment. In contrast, $\beta$-VAE, VAE and CausalVAE fail to achieve a good estimation of the true latent variables, since they are not identifiable. iVAE obtains unsatisfactory results, since its theory only holds for i.i.d. latent variables. Figure 8 shows the performance in recovering latent causal variables on synthetic data with different numbers of the latent causal variables. Figure 9 shows the structural Hamming distance (SHD), a standard metric for comparing graphs via their adjacency matrices, for the recovered skeletons (the undirected version of the graph with all directed edges replaced by undirected edges), causal graphs, and completed partial directed acyclic graphs (CPDAGs, which capture the set of all DAGs that are Markov equivalent), by SuaVE and CausalVAE. Since CausalVAE is not identifiable, it obtains fully connected graphs, the SHD of the recovered CPDAG is the same as one of the recovered skeletons.
@@ 7.1.1 Experiments on Changes of Part of Weights
In this part, we conduct experiments on changes of part of weighs, to verify the results in Theorem 4. To this end, we consider two cases on the following causal graph: $z_{1}\rightarrow z_{2}\rightarrow z_{3}\rightarrow z4$. The data details are similar to those of the 4-dimensional case mentioned in Appendix A.6.2, except that $z_{3}$ is a parent node of $z_{4}$. Case 1: the weight on the edge of $z_{1}\rightarrow z_{2}$ (and accordingly, $z_{2}\rightarrow z_{3}$, $z_{3}\rightarrow z_{4}$) remains unchanged across $\mathbf{u}$, while the weights on the remaining edges change across $\mathbf{u}$, Case 2: the weight on the edge of $z_{1}\rightarrow z_{2}$ (and accordingly, $z_{2}\rightarrow z_{3}$, $z_{3}\rightarrow z_{4}$) includes a constant part (i.e., $\lambda_{1,2}(\mathbf{u})+b$, $b$ is a non-zero constant.) across $\mathbf{u}$, while the weights on the remaining edges change across $\mathbf{u}$. Again, the network architecture, optimization, and hyper-parameters are the same as used for experiments of Figure 7 (a). Figures 10 and 11 show the performance of the proposed method in the settings corresponding to Case 1 and Case 2 mentioned above. According to MPC, we observe that unchanged weights in Case 1 or weights involving a constant term in Case 2 lead to non-identifiability. Conversely, changing weights contribute to the identifiability of the corresponding nodes. These empirical results align with the partial identifiability results in Theorem 4.
@@
Figure 11: Case 2: Performances of the proposed method with the change of part of weights. The ground truth of the causal graph is $z_{1}\rightarrow z_{2}\rightarrow z_{3}\rightarrow z_{4}$. From top to bottom: weight on $z_{1}\rightarrow z_{2}$ (and accordingly, $z_{2}\rightarrow z_{3}$, and $z_{3}\rightarrow z_{4}$) includes a non-zero constant part and changing part, i.e., $\lambda(\mathbf{u})+b$. Those results are consistent with the analysis of partial identifiability results in Theorem 4.
@@
Our identifiability result includes nonlinear ICA as a special case, in which the latent causal variables do not exhibit any causal relationships among themselves, effectively collapsing the causal structure into a fully disconnected graph. To verify this point, we conduct experiments on the case where any two latent causal variables $z_{i}$ are independent given $\mathbf{u}$. Data details are the same as mentioned in section A.6.2, but we here enforce $\lambda_{i,j}(\mathbf{u})=0$ for each $i,j$. Again, the network architecture, optimization and hyper-parameters are the same as section A.6.2, except for learning rate (here we set 1r=1e21r=1e-2). Figure 12 shows the performances of the proposed SuaVE and iVAE. We can see that iVAE is slightly better than the proposed SuaVE, both obtain satisfying results in terms of MPC.
@@ 7.1.3 Performance for violated assumption on distribution
Table 1: Performance for violated assumptions. Uniform (Laplace, Gamma) Noise: the distribution of latent noise variables. Matching Assumptions: a setting that matches our theoretical assumptions.
@@ 7.2 Image Data
We further verify our identifiability results and the proposed method on high-dimensional image data from the chemistry dataset proposed in Ke et al. (2021), which is corresponding to simple chemical reactions where the state of an element can cause changes to another variable’s state. The environment consists of a number of objects whose positions are kept fixed and thus, uniquely identifiable, while the colors (states) of the objects affected by this variable (according to the causal graph) can change. We use a weight-variant linear Gaussian model as mentioned in section A.6.2 for a 3-dimensional case, to generate the latent variables, and obtain the corresponding observational images. A visualization of the environment can be found in Figure 13.
@@
All these observations are consistent with the ground truth latent causal structure, i.e., $z_{1}\rightarrow z_{2}\rightarrow z_{3}$, together with results in Figure 14, demonstrating that the proposed method successfully learns latent causal structures. Similar, we inversigate the change in reconstructions while traversing each learned latent variable obtained by $\beta$-VAE (Higgins et al., 2017) and CausalVAE (Yang et al., 2021). Figure 16 and 17 show traversal results on the learned representations by $\beta$-VAE (Higgins et al., 2017) and CausalVAE (Yang et al., 2021), respectively. we observed that changing each learned variable results in changes to the colors of all objects. Together with the results in Figure 14, these suggests that $\beta$-VAE (Higgins et al., 2017) and CausalVAE (Yang et al., 2021) tend to learn representations that entangle the true underlying latent causal variables and fail to recover the true underlying latent causal structure.
@@
Following Ghassami et al. (2018), we applied the proposed method to fMRI hippocampus dataset (Laumann and Poldrack, 2015), which contains signals from six separate brain regions: perirhinal cortex (PRC), parahippocampal cortex (PHC), entorhinal cortex (ERC), subiculum (Sub), CA1, and CA3/Dentate Gyrus (DG) in the resting states on the same person in 84 successive days. Each day is considered as different $\mathbf{u}$, thus $\mathbf{u}$ is a 84-dimensional vector. Since we are interested in recovering latent causal variables, we treat the six signals as latent causal variables by applying a random nonlinear mapping on them to obtain observed data. We then apply various methods on the observed data to recover the six signals. It is worth noting that, with respect to Assumption (v), one of the key conditions in our theory, the context of fMRI might be interpreted as providing an approximate rather than a literal reference condition. For example, different cognitive tasks or resting states are known to substantially suppress or activate specific effective connections between brain regions. While such modulations may not constitute perfect hard interventions, they may weaken certain causal influences, serving as empirical analogues of the reference condition assumed in our theory. Our empirical results suggest that such approximate satisfaction of the assumption might already be sufficient in practice.
Figure 19 (a) shows the performance of the proposed SuaVE in comparison to iVAE, $\beta$-VAE, VAE and CausalVAE in recovering the latent six signals. $\beta$-VAE aims to recover independent latent variables, and it obtains an interesting result: enforcing independence (e.g., $\beta=25,50$) leads to worse MPC, and relaxing it (though contracting to its own independence assumption) improves the result (e.g., $\beta=4$). This is because the latent variables given the time index are not independent in this dataset.
@@
Figure 20 (a) show the edges in the anatomical ground truth reported in Ghassami et al. (2018). Figure 20 (b) show the latent recovered edges by the proposed SuaVE, in the setting where we treat fMRI data as latent causal variables. For comparison, Figure 20 (c) shows the recovered causal graph reported in (Ghassami et al., 2018), in the setting where fMRI data are treated as observed variables. That is, the setting in (b) is more challenging than the setting in (c).